OpenStudy (anonymous):

PLEASE HELP ME!! I WILL GIVE A MEDAL TO THE BEST ANSWER!!... Enter the answer. b + 6.7 = 19.4

OpenStudy (anonymous):

The easiest way to do this is to subtract 6.7 from 19.4. So, 19.4-6.7=12.7 Plug 12.7 in for B and voila :)

OpenStudy (anonymous):

b+6.7=19.4 12.7+6.7= 19.4

OpenStudy (anonymous):

Thanks a lot @BraceFacee !